Question 14 1 pts Monopoly Demand: Q 100 - 0.20P Cost: TC 10+ 50Q Solve for the profit-maximizing Price
Answer #1

Given

Q=100-0.20p

0.20p=100-Q

p=500-5Q

TR=p×Q=(500-5Q)Q=500Q-5Q^2

MR=d/dQ=500-10Q

TC=10+50Q

MC=d/dQ=50

In monpoly

at maximizing MR=MC

500-10Q=50

500-50=10Q

10Q=450

Q=45

Profit Maximizing price P=500-10(45)=500-450=$50